Astronomy Report – May 23th

Astronomy Report

Name:  Janet Biggs     Crew: 181
date: Sol 10 (23 May 2017)

Sky Conditions: Clear

Wind Conditions: Calm

Observation Start Time: 22:40

Observation End Time: 01:30

Summary: We finally got our two-star alignment!!!  We had a great night viewing and imaging.  Still lots to learn, but are totally excited about our first real steps!  Our entire crew came out and we took turns viewing and trying different exposures.

We used the Telerad, 9 x 50 finderscope,  Autoguide Scope, and laser pointer to find our two-star alignment.  Once we confirmed ALIGN SUCCES, we slewed to Jupiter.  Using the 25 mm eyepiece we were able to see three moons (As we are novices, please correct any incorrect identifications that we make).

We attached my Cannon 5D Mark III to the Autoguide Scope using the adapter Peter recommended.  We photographed Jupiter using a wide range of ISO settings and time exposures.  We then slewed to find M81.  Our two-star alignment may have been off (we only used three calibration stars), or it may have been our inexperience, or lack of magnification as we had my camera on the Autoguide rather than an eyepiece, but we never clearly identified M81 (although we got some interesting photos trying).  We then slewed to Saturn.  We were able to image Saturn, but struggled with a good focus on my camera.

We are eager to get back out there!

Info on attached images:

Jupiter, ISO 200, 5 sec.

Saturn, ISO 2000, 30 sec.

Attempt at M81, ISO 10000, 30 sec.

Objects Viewed: Jupiter, M81 (attempted), and Saturn

Problems Encountered: No problems, other than focusing my camera.


First image from the mission as we look for M81


Three of Jupiter’s moons



Astronomy Report – May 21st

Crew 181 Astronomy Report 21 May 2017

Name:  Janet Biggs     Crew: 181
Date: Sol 7 (20 May 2017)
Sky Conditions: Clear

Wind Conditions:

Observation Start Time: 22:30

Observation End Time: 24:30

Summary: Since I am a novice, the learning curve is a bit slower for me than anticipated.  Our Engineer Charlie has been working with me, which has been helpful.  We are both reading guides and manuals during the day to familiarizing ourselves with the equipment.  We are still trying to set our two-star alignment.  We were able to view and center Mizor last night in the finderscope.  Although it did not appear centered in the Telrad.

We are still having trouble focusing the telescopes 2” eyepiece.  If you can give us any tips here it would be appreciated as we cannot complete a two-star alignment until we master focusing.

I was able to take some video images of Charlie working inside the Observatory … one of my goals for my mission!

We are eager to try again, but last night was too cloudy and we had to get the station ready for our media guest.  Tonight, we have a media visit from Ozzy Ozbourne so I’m not sure if we will be able to try any viewing.

Objects Viewed: Mizar and Alcor

Problems Encountered: No problems, other than the above mentioned focusing problem.

Astronomy Report – April 17th

Hello Peter,

We tried yesterday night to focus on shining stars. We used the black focus first, then the gold one, we understand the way it is supposed to work, but here is the problem : the image is never sharp, and when we defocus the star become « twisted » in one direction. To illustrate the problem, I joined a sketch of our problem. The quality of the draw is awfull, I’m sorry, we can’t use internet so I did what I could. Mathieu thinks this problem is well known and could be explained by the misalignment of mirrors. Would you know what problem it is? What would it be caused to? And what could we do to fix it?

Best regards,

Quentin and Mathieu

Astronomy Report – April 16th

Astronomy Report

Name: Quentin Thomas, Mathieu Vander Donckt      Crew: 178
date: 14 April 2017

Sky Conditions: clear

Wind Conditions: windy, we closed the lower part of the shutter because it was shivering.

Observation Start Time: 1am

Observation End Time: 2am

Summary: The first goal was to set up the telescope for the first time. We couldn’t make any photography due to the problems described below.

Objects Viewed: Arcturus, Jupiter

Problems Encountered: impossible to focus. For each viewed star, what should have been a single point was stretched out. Jupiter was fuzzy. We guess this is due to the alignment of the primary mirror. Can you send us the procedure to follow? Thank you.

Also, we didn’t dare to open the lower part of the shutter because of the wind. It was shivering. We let the upper part open to set up the telescope for next nights. After encountering the focusing problem, we closed everything.


Best regards,

Quentin “Dips” Thomas, Crew Astronomer,

Mathieu “Mitch” Vander Donckt, Crew Scientist and Journalist.


Geology Report – March 30th

rew 177 Geology Report  March 30, 2017
Crew 177 Geologist Report
Dear Mission Support,
Today is a sunny day, and we are ready to head to the geology stop “The Little Canyon”. After the delicious pancake breakfast, Becky, Esteban, Pitchayaba, and I went on the two Rovers and one ATVs with the measure tools – rope and tape.
Our mission for today’s EVA is to measure the height of the little canyon. About 15 mins, we have arrived the little canyon, and I climbed to the top of the canyon and throw the rope down to the cliff, but I held the one end of the rope, and then Esteban made a mark on the rope. We took some great pictures and videos over there. After we came back to the HAB, we measured the length of the rope, the height of the little canyon is around 18 feet. It was a great experience.
Caleb Li

Astronomy Report – March 30th

Crew 177 Astronomer Report 30Mar2017
Astronomer Report
On Sol 1 & 2 the sky was too cloudy to attempt to use the telescope.  On Sol 3, we enjoyed the night sky with the naked eye.  On Sol 4 we went to the Musk Observatory and followed the quick guide procedures.
We were able to view and focus on Jupiter, but were unable to focus on Polaris or Beetle Juice.  Because we were unable to complete the alignment procedure we discontinued use of the telescope.
Becky Parker
Crew 177 Astronomer

Geologist & HSO Report – March 28th

Crew 177 Geologist & HSO Officer


This is the third day being this HAB. I still cannot get rid of the excited feeling. Today is also the media day after we woke up in the morning around 7 am, we had the interview with the National Public Radio journalist, Ms. Rae Bichell. In the morning, four people from our crew and the reporter went out for a short EVA at the Cow Patty Field. In the afternoon, Dr. Villarroel, Joseph Quass, Victoria LaBarre and I went out to the Geology EVA. We ride 2 ATVs and 1 Rover, we went to the two stops, first stop is Cow Patty field, and we collected white-coating rock sample from the geology site 1A and 1B, sketched the sedimentary structure and took some photos of the field. We spent 50 mins on this stop. The next stop is the Little Canyon, we measured the height of the canyon, collected 8 bags of the dirt samples at the canyon, and took some photos there. We spent around 50 mins. Then, we came back to the HAB to enjoy the delicious tortilla meal.


Crew health situation: Everyone feels good, but we are worried about the lack of toilet and it might affect health.


Caleb Li

Science Report – March 27th

Science Report
Crew 177
Filled by Esteban Ramirez, Elijah Espinoza, Victoria LaBarre, Joseph Quaas.

Today, crew 177 spent the day working on a Chemistry EVA as well as working on their independent projects.  Unfortunately one of the goal of the chemistry EVA could be completed due technical difficulty with the Geiger counter.  Crew member will work on trying to solve the problem as a secong group will be send later this week to complete the task. Next, crew went to the Candor Chasma to collect water samples. Water was not located in this area although several moist areas were identified under rocks. Gypsum was evaluated and the students discussed that you can make water from gypsum if needed.  Some samples were brought back for future composition evaluation.  An important amount of green vegetation and colorful were observed down in the canyon.  We observed birds, and found evidence of nesting by some unknown creatures.

Esteban Espinoza, spent part of the afternoon working on his project. The generator bike was unpacked and partially assembled during SOL 1. During SOL 2, the final components to the bike were attached and we began test trials. Crew member, Pitchayapa Jingjit began the first test trial. There were a few complications with the readings from her heartrate sensor initially, but these were later solved. The bike generator was also initially unable to generate power due to weak initially battery levels, but this issue was solved using a larger battery to charge the generator’s batteries. In Pitchayapa’s second trials, we could obtain constant heartrate measurements, and there was a significant increase in the battery’s voltage. We plan to begin official trials in the coming days wherever possible.
Caleb Li spent part of his afternoon working on  a LED digital clock in the air lock, since there is not a clock in the air lock so far; he came up the idea that make a digital clock by using FPGAs and the Basys 2 board. This clock could optimize the crew member’s experience when they wait in the air lock and go out do the simulation. He is currently working on the alarm function on the Basys 2 board, he found some issues related with the alarm codes and he is working on solve them. He is planning to continue troubleshooting for the alarm function tomorrow. Hope it will work out.

Joseph Quass, in addition to led the chemistry’s EVA during the morning, he spent his afternoon working on a virtual reality training simulator. This specific simulation is based on MDRS, so the sim would train incoming crews of the basic layout of the HAB, as well as important surrounding locations. Today, he worked on editing textures in order to fix some clipping issues that were occurring prior to my arrival at MDRS. The rest of his time at MDRS will be devoted to making the simulation as accurate as possible to the real deal in order to immerse users. The goal of the simulation is to train participants in the layout and through certain situations, such as rescue emergencies, that can occur on-site. Tomorrow, he will continue with bug-fixes and terrain creation.

Elijah and Victoria are doing joint research about robotic. First the robot was unpacked and parts were checked for damage. They proceeded to assemble the robot, starting with the frame, circuitry, wheels and then finally the motors and gearboxes. Once fully assembled, Victoria taught Elijah how to use a soldering iron so that we could connect motor wires to motor controller wires, which went into the brain. This ended up becoming larger tutorial for other members of the crew who were in tested in learning how to solder wire. After connecting the wires, they switched on the robot. They noted that although the robot turned on, the wireless communication between the remote control and the brain seemed to have trouble connecting.  They plan to try solving this problem by cutting up two USB cables that they brought to solder together a USB to USB cable that will connect the brain and the remote control.

Geologist Report – March 26th

Crew 177 Geologist & HSO Report

My name is Caleb Li and my roles in Crew 177 are the geologist and Health Safety Officer. Today is our first day to have EVA, and we did EVA orientation with four crew members in the morning and the other four members in the afternoon. We learned how to operate the radio and GPS connection, drive ATVs safely, observe the variation of the rock and landscape, collect the rock samples, and take pictures. It is fun and very interesting to put on the space suit and experience the astronaut’s feelings.

Caleb Li

Science Report – March 25th

Science Report

Michał Kazaniecki
Crew 176 Engineer

In recent days we had 2 EVAs with Ares rover. We performed some field tests to check procedures and collaboration with rover and to examine Ares behavior in various conditions.

Two days ago we had objectives regarding human-robot interactions. Ares was helping geologist with collecting the samples. It took some pieces of rock and delivered it to destination.

Yesterday we performed further field tests. We headed north, to the area between Hab and North Ridge. Ares climbed on several quite steep hills, some pf them up to 35 degrees. It was a great opportunity to observe how the suspension works when forces on each wheel are not equal (and asymmetrical). Besides, we checked the performance of motors and grip of wheels on loose soil as well as we drove in the dust and mud to verify the impact of water or dirt on mechanical parts.

One of unexpected events was when Ares dug itself in sand. Unfortunately, it blew two electric fuses while it was trying to escape.

We also performed tests on maximum range of wireless communication (without additional antenna). Results were quite surprising – in open field (without any obstacles around) we nearly lost visual contact with rover but the connection was still quite good.

That EVA was also an opportunity to train an additional rover operator. It took less than minute to explain how to drive a rover.

A completely new experience, impossible to gain on robotic competitions like University Rover Challenge  was operating the rover in helmet and space suit. We found out, that using a joystick in gloves is as comfortable as without them.  On the contrary, sun reflections on the helmet makes nearly impossible to see anything on computer screen.